# From simplex slicing to sharp reverse Hölder inequalities

### v1

@@ Abstract.
2020 Mathematics Subject Classification. Primary 52A40; Secondary 52B12.
@@ 1.1. Motivation
The problem of finding minimal and maximal volume sections — critical sections — of various convex bodies has received significant attention over the last several decades. This topic, being somewhat tangential to the broad area of geometric tomography, has nevertheless played a pivotal role in its development, as witnessed for instance by a strikingly simple counter-example to the Busemann-Petty problem thanks to Ball’s famed cube slicing result (see [2, 3] for details).
Notably, the study of critical sections has also been fruitful in developing robust methods, with Fourier analytic ideas at the core (see [25]), only recently to be souped-up by a nascent probabilistic point of view involving negative moments of weighted sums of random variables (see, e.g. [11, 12, 15]). We refer to the recent survey [32] showcasing this approach, as well as contextualising it further, with comprehensive references and historical account. This very line of thought serves as the main motivation for this paper.
Our starting point is Webb’s simplex slicing result from [36] and the question whether it admits a probabilistic extension to negative moments, akin the aforementioned Ball’s cube slicing generalised in such a way in [11].
@@
$\Delta_{n}\coloneq\left\{x\in\mathbb{R}^{n+1},\ \sum_{j=1}^{n+1}x_{j}=1,x_{1},\dots,x_{n+1}\geq 0\right\}$
embedded in the hyperplane :-$\mathcal{H}\coloneq\{x\in\mathbb{R}^{n+1},\ \sum x_{j}=1\}$ in $\mathbb{R}^{n+1}$ (that is, the vertices of $\Delta_{n}$ are the standard basis vectors $e_{1},\dots,e_{n+1}$).
@@
$\max_{H}\ \operatorname{vol}_{n-1}(\Delta_{n}\cap H)=\frac{\sqrt{n+1}}{(n-1)!}\cdot\frac{1}{\sqrt{2}},$
where the maximum is taken over all affine subspaces $H$ of $\mathcal{H}$ of (relative) codimension $1$ constrained to pass through the barycentre $\frac{1}{n+1}(1,\dots,1)$ of $\Delta_{n}$. Plainly, every such hyperplane $H$ extends to a codimension $1$ subspace of $\mathbb{R}^{n+1}$ by taking the affine hull of $H$ and the origin, yielding a hyperplane $a^{\perp}\coloneq\{x\in\mathbb{R}^{n+1},\ \left\langle a,x\right\rangle=0\}$ in $\mathbb{R}^{n+1}$ with an outer-normal vector $a=(a_{1},\dots,a_{n+1})$ satisfying $\sum_{j=1}^{n+1}a_{j}=0$ (so that $a^{\perp}$ contains the barycentre of $\Delta_{n}$). (We use the standard Euclidean structure given by the inner product $\left\langle x,y\right\rangle\coloneq\sum x_{j}y_{j}$ and resulting Euclidean norm $|x|\coloneq\sqrt{\left\langle x,x\right\rangle}$ for arbitrary vectors $x$ and $y$.) With this identification and the normalisation $|a|=1$, we have the following formula for volume of sections, instrumental in Webb’s work,
@@
For completeness, we include a standard derivation in the appendix. Here $f_{\sum a_{j}\mathcal{E}_{j}}$ is (the continuous version of) the density of the random variable $\sum_{j=1}^{n+1}a_{j}\mathcal{E}_{j}$, and $\mathcal{E}_{1},\dots,\mathcal{E}_{n+1}$ are independent identically distributed (i.i.d.) standard exponential random variables (that is with density $e^{-x}\textbf{1}_{(0,+\infty)}(x)$ on $\mathbb{R}$). Webb’s result thus amounts to the following sharp upper bound on the density at 00 of weighted sums of i.i.d. exponentials $\mathcal{E}_{j}$,
@@
At the heart of the aforementioned negative moments paradigm lies the following elementary fact: for an integrable function $f$ on $\mathbb{R}$ which is, say, continuous at 00, we have
@@
for instance see Lemma 4 in [20] as well as Lemma 4.3 in [12] for a multivariate version. In view of this observation, Webb’s result can be restated in a yet another equivalent form, as the bound
@@
for all unit vectors $a$ with $\sum a_{j}=0$. Does this bound continue to hold without taking the limit, say for all fixed $p$ in some neighbourhood of $-1$?
@@
Recall that a real-valued random variable $X$ is called log-concave, if it is continuous with a log-concave density $f$ on $\mathbb{R}$, that is a function of the form $f=e^{-\phi}$ with a convex function $\phi\colon\mathbb{R}\to(-\infty,+\infty]$. For instance, a uniform random variable on an interval, or a Gaussian random variable is log-concave, and most importantly for this discussion, exponential random variables are log-concave. Moreover, it is well-known that convolutions of log-concave functions are log-concave (by the Prékopa-Leindler inequality), thus sums of independent log-concave random variables are log-concave. They naturally play a prominent role in modern convex geometry, see the monographs [1, 8] for background. Needless to say, the weighted sums
@@
of independent exponential random variables $\mathcal{E}_{j}$ have naturally arose in simplex slicing, as just considered. These are log-concave random variables. Owing to the geometric constraints imposed on the weights $a_{j}$, they are centred, that is with mean 00,
@@
For every $-1<p\leq 1$ and every mean 00 log-concave random variable $X$, we have
@@
$\|X\|_{p}\coloneq(\mathbb{E}|X|^{p})^{1/p}\ \in[0,+\infty],$
with the usual convention of adopting the limiting expressions as definitions at $p=0$ and $p=+\infty$, respectively as $\|X\|_{0}\coloneq e^{\mathbb{E}\log|X|}$ (the geometric mean) and $\|X\|_{\infty}\coloneq\text{ess sup}|X|$ (the essential supremum). For log-concave random variables $X$, $\|X\|_{p}$ is finite for all $p\in(-1,+\infty)$ as a result of an exponential decay of their densities (see, e.g. Lemma 2.2.1 in [8]).
In particular, for any log-concave random variable $X$ with mean 00 variance $1$, and for every $-1<p<0$, Theorem 1 yields
@@
Theorem 1 is in fact obtained as a corollary to a sharp $L_{p}-L_{1}$ moment comparison inequality, where in lieu of the variance constraint, we impose the $L_{1}$ constraint (which goes hand in hand with the mean 00 constraint, but more on that later). Intriguingly, there is a phase transition of the extremising distribution. Specifically, for $p\geq 1$, we define
(4) $C_{p}\coloneq\max\left\{\Gamma(p+1)^{1/p},\ \frac{e}{2}\|\mathcal{E}-1\|_{p}\right\},$
@@
For every log-concave random variable $X$ with mean 00 and every $-1<p\leq 1$, we have
@@
Theorem 1 follows at once. Indeed, using Theorem 2 twice, for $-1<p\leq 1$, we have
@@
For every log-concave random variable $X$ with mean 00 and every $-1<p\leq 1\leq q\leq p_{0}$, we have
@@
Log-concave random variables constitute a prominent class enjoying Khinchin-type inequalities, with natural connections to convex geometry (marginals of uniform measures on convex bodies are log-concave), which can be traced back at least to the seminar paper [28] by Milman and Pajor; we also refer to Chapter 5 in the lecture notes [19]. Even preceding this development (arising in a context of distributions with monotone hazard rate significant in statistics with applications to renewal processes), notable is the result of [5] where (8) was established for all $-1<p<q$ for random variables $X$ which are symmetric or nonnegative (a.s.) with log-concave tails, i.e. when $t\mapsto-\log\mathbb{P}\left(X>t\right)$ is convex (implied when $X$ is log-concave). For Khinchin-type inequalities like (8) with sharp constants, weakening the symmetry assumption however has proved challenging and has been investigated only recently (in the case of actual $L_{p}$ norms). Given $1\leq p\leq q$, let $C_{p,q}^{*}$ be the best constant $C$ such that $\|X\|_{q}\leq C\|X\|_{p}$ holds for all log-concave random variables with mean 00. Eitan in [16] (Theorem 3.1 therein) has showed that $C_{p,q}^{*}=\max_{0\leq s\leq 1}\frac{\|Z_{s}\|_{q}}{\|Z_{s}\|_{p}}$, that is $C_{p,q}^{*}$ is attained within the one parameter family of two-sided exponential random variables $Z_{s}\coloneq s(\mathcal{E}-1)-(1-s)(\mathcal{E}^{\prime}-1)$, $0\leq s\leq 1$, $\mathcal{E},\mathcal{E}^{\prime}$ being i.i.d. standard exponential random variables. Eitan has also conjectured that when $p,q$ are even integers, the maximum is attained at $s=0,1$ (a one-sided exponential), and confirmed this for $p=2$, and arbitrary even $q$, as well as for all even integers $p\leq q\leq 100$. Nota bene, he has provided nice applications to geometry of convex bodies characterising simplices as bodies with heaviest tails. Murawski in [29] has established a concrete clean bound $C_{p,q}^{*}\leq\frac{p}{q}$.
Our main result completely determines $C_{1,q}^{*}$ (as well as $C_{p,1}^{*}$ when $-1<p\leq 1$), revealing a phase transition of the extremising distribution.
@@ 2.1. Overall strategy and the main idea: $L_{1}$ proxy
where the $\inf\big{/}\sup$ (depending on the sign of $p$) is taken over all log-concave functions $f\colon\mathbb{R}\to[0,+\infty)$. The objective functional is in fact linear in $f$, so are the constraints. After a compactification of the domain, this naturally lands itself into a problem where the so-called localisation techniques developed in [17] may be applied. As such (skipping many details), the problem is then reduced to optimisation over the log-concave functions on a compact interval with at most $3$ degrees of freedom, leading to an explicit but complicated optimisation problem with $4$ parameters (perhaps intractable, without any further nontrivial ideas).
@@
Crucially, the mean 00 constraint $\int_{\mathbb{R}}xf(x)\mathrm{d}x=0$ is plainly equivalent to
@@
are fixed to be $\frac{1}{2}$. Together with the convexity/concavity of $x\mapsto|x|^{p}$ on $(-\infty,0)$ and $(0,+\infty)$, this is then leveraged in Step I by a simple optimisation argument on each of the half-lines $(-\infty,0)$ and $(0,+\infty)$ (without the need of compactification), reducing the whole problem to densities $f=e^{-\phi}$ with piece-wise linear potentials $\phi$ having only two slopes, resulting in an explicit opitmisation problem over only $1$ parameter, then tackled in Step II.
@@
The workhorse of the main reduction in Step I is a simple but quite powerful method of proving integral inequalities by analysing the number of sign changes of the integrand. It will also be instrumental for many tweaks in Step II to keep painstaking technicalities to a minimum. We shall illustrate this technique now with a toy example (needed later anyway). Let $0<x_{1}<x_{2}<\dots<x_{k}$. We say that function $f\colon(0,+\infty)\to\mathbb{R}$ has sign changes at $x_{1},x_{2},\dots,x_{k}$ with sign pattern (+,,)(+,-$\cdots$)( + , - , ⋯ ) (resp. (,+,)(-,+$\cdots$)( - , + , ⋯ )) if $f$ is positive (resp. negative) on $(0,x_{1})\cup(x_{2},x_{3})\cup\cdots$ and $f$ is negative (resp. positive) on $(x_{1},x_{2})\cup(x_{3},x_{4})\cup\cdots$. Similarly, we say that $f$ has weak sign changes at $x_{1},x_{2},\dots,x_{k}$ with sign pattern $(0+,0-,\cdots)$ (resp. $(0-,0+,\cdots)$) if $f$ is nonnegative (resp. nonpositive) on $(0,x_{1})\cup(x_{2},x_{3})\cup\cdots$ and $f$ is nonpositive (resp. nonnegative) on $(x_{1},x_{2})\cup(x_{3},x_{4})\cup\cdots$. If there exist $0<x_{1}<x_{2}<\dots<x_{k}$ such that $f$ has (weak) sign changes at $x_{1},x_{2},\dots,x_{k}$, we say that $f$ admits a finite (weak) sign change pattern.
@@
It is not true that for every smooth function $f\colon(0,+\infty)\to\mathbb{R}$ and every finite interval $[a,b]\subset(0,+\infty)$, if $f$ takes both positive and negative values on $[a,b]$ then $f\big{|}_{[a,b]}$ admits a finite weak sign change pattern. However, it becomes true if “smooth” is replaced with “analytic” or “convex”.
@@
Let $h\colon(0,+\infty)\to\mathbb{R}$ be an integrable function. Suppose $\int_{0}^{+\infty}h(x)\mathrm{d}x=0$. If $h$ has exactly one weak sign change with sign pattern $(0-,0+)$, and if $h$ is nonzero on a set of positive measure, then
@@
By the constraint that $h$ integrates to 00, we have for every $\alpha$,
@@
This concept appears naturally in moment problems, was advanced in [14], and has been extensively used in a variety of contexts, see [6, 7, 10, 16, 22, 26, 27, 31]. Robust tools of this sort can be traced back to the classical work [23] of Karlin and Novikoff (with a slightly different focus on characterising certain cones of functions and their duals).
@@
(9) $X_{a,b}\coloneq a(\mathcal{E}-1)-b(\mathcal{E}^{\prime}-1).$
@@
(with the convention that for $a=0$, or $b=0$, the first, or the second case is empty, respectively, where we put $g_{a,b}\equiv 0$). In words, the distribution of $X_{a,b}$ is a mixture of two one-sided exponentials: $a\mathcal{E}-a+b$ with probability $\frac{a}{a+b}$ and $-b\mathcal{E}^{\prime}-a+b$ with probability $\frac{b}{a+b}$.
@@
Let $X$ be a log-concave random variable with mean 00. There are unique $a,b\geq 0$ such that
@@
Let $\alpha\coloneq\mathbb{P}\left(X>0\right)$. Without loss of generality, we can assume that $\alpha\leq\frac{1}{2}$ (otherwise, we consider $-X$ instead of $X$). Since $\mathbb{E}X=0$, by Grünbaum’s theorem ([8, Lemma 2.2.6]), in fact, $\alpha\in[\frac{1}{e},\frac{1}{2}]$. Suppose that $a\geq b$. Then,
@@
Since the function $u\mapsto\frac{1}{e}\frac{e^{u}}{1+u}$ is continuous and strictly increasing on $[0,{+\infty})$, it is bijective between $[0,1]$ and $[\frac{1}{e},\frac{1}{2}]$. As a result, there is a unique value of the ratio $b/a$ for which $\mathbb{P}\left(X_{a,b}>0\right)=\alpha$. Furthermore, the mean of $X_{a,b}$ is 00, so
@@
{asy}
unitsize(1.7cm); draw((-3,0)–(3,0),arrow=Arrow()); draw((0,-1)–(0,3),arrow=Arrow()); draw((-2,-1)–(-1,2.7)–(2.5,-1),p=linewidth(2)); label(”$\log g$”,(2.8,-1)); draw((-1,-0.1)–(-1,0.1)); label(”$m$”,(-1,-0.3)); draw((-1.9,-1)–(-1.8,-0.2)–(-1.7,0.5)–(-1.6,1.1)–(-1.5,1.6)–(-1.4,2.0)–(-1.3,2.3)–(-1.2,2.37)–(-1.1,2.36)–(-1.0,2.34)–(-0.5,2)–(0.8,0.9)–(1,0.7)–(1.2,0.45)–(1.3,0.3)–(1.4,0.1)–(1.7,-0.4)–(2,-1),p=dotted); label(”$\log f$”,(1.6,-1)); draw(circle((-1.82,-0.33),0.04)); draw(circle((-1.09,2.36),0.04)); draw(circle((0.3,1.32),0.04)); draw(circle((1.33,0.23),0.04));
@@
We proceed with showing (12). By the homogeneity of $X/\mathbb{E}|X|$, we can assume that $\mathbb{E}|X|=1$. Let $f\colon\mathbb{R}\to[0,{+\infty})$ denote the density of $X$. For brevity, we shall write $g=g_{a,b}$ for the density of $X_{a,b}$ given above. These are log-concave probability densities, by construction $\log g$ is piecewise linear with maximum at $m\coloneq-(a-b)\leq 0$, enjoying the constraints
@@
Note that $f$ and $g$ on $[m,{+\infty})$ can intersect at most twice, that is $g-f$ changes sign at most twice because $\log f$ is concave and $\log g$ is linear. Similarly on $(-\infty,m]$. Because of the integral constraints on $[0,{+\infty})$, Lemma 5 asserts that $g-f$ needs to change its sign at least twice on $[0,{+\infty})$, say at $0\leq x_{1}<x_{2}$, unless $f=g$ (identically). Then, $g-f$ has the sign pattern (+,,+)(+,-,+)( + , - , + ) on $(0,{+\infty})$. Incidentally (it will be needed in the sequel), we also remark that $g-f\geq 0$ on $[m,0]$ as $\log g$ is linear on the entire half-line $[m,{+\infty})$. We take $\lambda,\mu\in\mathbb{R}$ such that $\psi(x)-\lambda x-\mu$ vanishes at $x_{1},x_{2}$ ($\lambda,\mu$ are thus uniquely determined by the corresponding linear equations). Since $\psi$ is convex, $\psi(x)-\lambda x-\mu$ enjoys the sign pattern $(0+,0-,0+)$. Then,

### v2

@@ Abstract.
Dedicated to Keith Ball, on his 65th birthday.
2020 Mathematics Subject Classification. Primary 52A40; Secondary 60E15, 52B12.
@@ 1.1. Motivation
The problem of finding minimal and maximal volume sections — critical sections — of various convex bodies has received significant attention over the last several decades. This topic, being somewhat tangential to the broad area of geometric tomography, has nevertheless played a pivotal role in its development, as witnessed for instance by a strikingly simple counterexample to the Busemann-Petty problem thanks to Ball’s famed cube slicing result (see [2, 3] for details).
Notably, the study of critical sections has also been fruitful in developing robust methods, with Fourier analytic ideas at the core (see [27]), only recently to be enhanced by a nascent probabilistic point of view involving negative moments of weighted sums of random variables (see, e.g. [12, 13, 16]). We refer to the recent survey [38] showcasing this approach, as well as contextualising it further, with comprehensive references and historical account. This very line of thought serves as the main motivation for this paper.
Our starting point is Webb’s simplex slicing result from [41] and the question whether it admits a probabilistic extension to negative moments, akin to the aforementioned Ball’s cube slicing generalised in such a way in [12].
@@
$\Delta_{n}=\left\{x\in\mathbb{R}^{n+1},\ \sum_{j=1}^{n+1}x_{j}=1,\ x_{1},\dots,x_{n+1}\geq 0\right\}$
embedded in the hyperplane $\mathcal{H}=\{x\in\mathbb{R}^{n+1},\ \sum x_{j}=1\}$ in $\mathbb{R}^{n+1}$ (that is, the vertices of $\Delta_{n}$ are the standard basis vectors $e_{1},\dots,e_{n+1}$).
@@
$\max_{{L}}\ \operatorname{vol}_{n-1}(\Delta_{n}\cap H)=\frac{\sqrt{n+1}}{(n-1)!}\cdot\frac{1}{\sqrt{2}},$
where the maximum is taken over all affine subspaces $L$ of $\mathcal{H}$ of (relative) codimension $1$ constrained to pass through the barycentre $\frac{1}{n+1}(1,\dots,1)$ of $\Delta_{n}$. Plainly, every such hyperplane $L$ extends to a codimension $1$ subspace of $\mathbb{R}^{n+1}$ by taking the affine hull of $L$ and the origin, yielding a hyperplane $a^{\perp}=\{x\in\mathbb{R}^{n+1},\ \left\langle a,x\right\rangle=0\}$ in $\mathbb{R}^{n+1}$ with an outer-normal vector $a=(a_{1},\dots,a_{n+1})$ satisfying $\sum_{j=1}^{n+1}a_{j}=0$ (so that $a^{\perp}$ contains the barycentre of $\Delta_{n}$). (We use the standard Euclidean structure given by the inner product $\left\langle x,y\right\rangle=\sum x_{j}y_{j}$ and resulting Euclidean norm $|x|=\sqrt{\left\langle x,x\right\rangle}$ for arbitrary vectors $x$ and $y$.) With this identification and the normalisation $|a|=1$, we have the following formula for volume of sections, instrumental in Webb’s work,
@@
For completeness, we include a standard derivation in the appendix. Here $f_{\sum a_{j}\mathcal{E}_{j}}$ is (the continuous version of) the density of the random variable $\sum_{j=1}^{n+1}a_{j}\mathcal{E}_{j}$, and $\mathcal{E}_{1},\dots,\mathcal{E}_{n+1}$ are independent identically distributed (i.i.d.) standard exponential random variables (that is with density $e^{-x}\textbf{1}_{(0,+\infty)}(x)$ on $\mathbb{R}$). Webb’s result thus amounts to the following sharp upper bound on the density at 0 of weighted sums of i.i.d. exponentials $\mathcal{E}_{j}$,
@@
At the heart of the aforementioned negative moments paradigm lies the following elementary fact: for an integrable function $f$ on $\mathbb{R}$ which is, say, continuous at 0, we have
@@
for instance see Lemma 4 in [22] as well as Lemma 4.3 in [13] for a multivariate version. In view of this observation, Webb’s result can be restated in yet another equivalent form, as the bound
@@
for all unit vectors $a$ with $\sum a_{j}=0$. Does this bound continue to hold without taking the limit, say for all fixed $p$ in some neighbourhood of 1-1?
@@
[…]
@@
of independent exponential random variables $\mathcal{E}_{j}$ have naturally arisen in simplex slicing, as just considered. These are log-concave random variables. Owing to the geometric constraints imposed on the weights $a_{j}$, they are centred, that is, with mean 0,
@@
For every $1<p\leq 1$ and every mean 0 log-concave random variable $X$, we have
@@
$\|X\|_{p}=(\mathbb{E}|X|^{p})^{1/p}\ \in[0,+\infty],$
with the usual convention of adopting the limiting expressions as definitions at $p=0$ and $p=+\infty$, respectively as $\|X\|_{0}=e^{\mathbb{E}\log|X|}$ (the geometric mean) and $\|X\|_{\infty}=\text{ess sup}|X|$ (the essential supremum). For log-concave random variables $X$, $\|X\|_{p}$ is finite for all $p\in(-1,+\infty)$ as a result of an exponential decay of their densities (see, e.g. Lemma 2.2.1 in [8]).
In particular, for any log-concave random variable $X$ with mean 0 and variance $1$, and for every 1<p<0-1<p<0, Theorem 1 yields
@@
Theorem 1 is in fact obtained as a corollary to a sharp $L_{p}-L_{1}$ moment comparison inequality, where in lieu of the variance constraint, we impose the $L_{1}$ constraint (which goes hand in hand with the mean 0 constraint, but more on that later). Intriguingly, there is a phase transition of the extremising distribution. Specifically, for $p\geq 1$, we define
(4) $C_{p}=\max\left\{\Gamma(p+1)^{1/p},\ \frac{e}{2}\|\mathcal{E}-1\|_{p}\right\},$
@@
For every log-concave random variable $X$ with mean 0 and every $1<p\leq 1$, we have
@@
Theorem 1 follows at once. Indeed, using Theorem 2 twice, for $1<p\leq 1$, we have
@@
For every log-concave random variable $X$ with mean 0 and every $1<p\leq 1\leq q\leq p_{0}$, we have
@@
Log-concave random variables constitute a prominent class enjoying Khinchin-type inequalities, with natural connections to convex geometry (marginals of uniform measures on convex bodies are log-concave), which can be traced back at least to the seminal paper [34] by Milman and Pajor; we also refer to Chapter 5 in the lecture notes [21]. Even preceding this development (arising in a context of distributions with monotone hazard rate significant in statistics with applications to renewal processes), notable is the result of [5] where (8) was established for all 1<p<q-1<p<q for random variables $X$ which are symmetric or nonnegative (a.s.) with log-concave tails, i.e. when $t\mapsto-\log\mathbb{P}\left(X>t\right)$ is convex (implied when $X$ is log-concave). For Khinchin-type inequalities like (8) with sharp constants, weakening the symmetry assumption however has proved challenging and has been investigated only recently (in the case of actual $L_{p}$ norms). Given $1\leq p\leq q$, let $C_{p,q}^{*}$ be the best constant $C$ such that $\|X\|_{q}\leq C\|X\|_{p}$ holds for all log-concave random variables with mean 0. Eitan in [17] (Theorem 3.1 therein) has shown that $C_{p,q}^{*}=\max_{0\leq s\leq 1}\frac{\|Z_{s}\|_{q}}{\|Z_{s}\|_{p}}$, that is $C_{p,q}^{*}$ is attained within the one parameter family of two-sided exponential random variables $Z_{s}=s(\mathcal{E}-1)-(1-s)(\mathcal{E}^{\prime}-1)$, $0\leq s\leq 1$, $\mathcal{E},\mathcal{E}^{\prime}$ being i.i.d. standard exponential random variables. Eitan has also conjectured that when $p,q$ are even integers, the maximum is attained at $s=0,1$ (a one-sided exponential), and confirmed this for $p=2$, and arbitrary even $q$, as well as for all even integers $p\leq q\leq 100$. Nota bene, he has provided nice applications to geometry of convex bodies characterising simplices as bodies with heaviest tails. Murawski in [35] has established a concrete clean bound $C_{p,q}^{*}\leq\frac{p}{q}$.
Our main result completely determines $C_{1,q}^{*}$ (as well as $C_{p,1}^{*}$ when $1<p\leq 1$), revealing a phase transition of the extremising distribution.
It is also worth mentioning a standard application to maximal volume central sections of isotropic convex bodies: the limiting case $p\searrow-1$ of Theorem 2 recovers Fradelizi’s result [19] (Corollary 1 therein), which is in the spirit of a main result of [33] (Corollary 4 therein establishes a sharp lower bound on noncentral sections).
@@ 2.1. Overall strategy and the main idea: $L_{1}$ proxy
where the $\inf\big/\sup$ (depending on the sign of $p$) is taken over all log-concave functions $f\colon\mathbb{R}\to[0,+\infty)$. The objective functional is in fact linear in $f$, so are the constraints. After a compactification of the domain, this naturally lands itself into a problem where the so-called localisation techniques developed in [20] may be applied. That is (skipping many details), through the Krein-Milman theorem and the identification of extreme points as log-concave functions on a compact interval with at most $3$ degrees of freedom (in the sense of [20]), the above optimisation problem is reduced to the family of compactly supported log-concave functions $f=e^{-\phi}$ with convex potentials $\phi$ which are piece-wise affine with at most two linear pieces, leading to an explicit but complicated optimisation problem with $4$ parameters (perhaps intractable, without any further nontrivial ideas). See, e.g. [7, 28, 30, 35] for further details, where this exact scheme has been employed. Depending on how one sets the parameters up for the reduced problem, this might amount to, say, given $a,b>0$, extremising $\int_{-a}^{b}|x|^{p}f(x)\mathrm{d}x$ with $f(x)=c\exp(-\max\{\alpha x+\beta,\alpha^{\prime}x+\beta^{\prime}\})$ with the parameters $a,b,c,\alpha,\alpha^{\prime},\beta,\beta^{\prime}$ constrained by $\int f=1,\int xf(x)=0$, $\int x^{2}f(x)=1$, leaving out $4$ free parameters.
@@
Crucially, the mean 0 constraint $\int_{\mathbb{R}}xf(x)\mathrm{d}x=0$ is plainly equivalent to
@@
are fixed to be $\frac{1}{2}$. Together with the convexity/concavity of $x\mapsto|x|^{p}$ on $(-\infty,0)$ and $(0,+\infty)$, this is then leveraged in Step I by a simple optimisation argument on each of the half-lines $(-\infty,0)$ and $(0,+\infty)$ (without the need of compactification), reducing the whole problem to densities $f=e^{-\phi}$ with piece-wise linear potentials $\phi$ having only two slopes, resulting in an explicit optimisation problem over only $1$ parameter, then tackled in Step II.
@@
The workhorse of the main reduction in Step I is a simple but quite powerful method of proving integral inequalities by analysing the number of sign changes of the integrand. It will also be instrumental for many tweaks in Step II to keep painstaking technicalities to a minimum. We shall illustrate this technique now with a toy example (needed later anyway). Let $0<x_{1}<x_{2}<\dots<x_{k}$. We say that function $f\colon(0,+\infty)\to\mathbb{R}$ has sign changes at $x_{1},x_{2},\dots,x_{k}$ with sign pattern $(+,-,\cdots)$ (resp. $(-,+,\cdots)$) if $f$ is positive (resp. negative) on $(0,x_{1})\cup(x_{2},x_{3})\cup\cdots$ and $f$ is negative (resp. positive) on $(x_{1},x_{2})\cup(x_{3},x_{4})\cup\cdots$. Similarly, we say that $f$ has weak sign changes at $x_{1},x_{2},\dots,x_{k}$ with sign pattern $(0+,0-,\cdots)$ (resp. $(0-,0+,\cdots)$) if $f$ is nonnegative (resp. nonpositive) on $(0,x_{1})\cup(x_{2},x_{3})\cup\cdots$ and $f$ is nonpositive (resp. nonnegative) on $(x_{1},x_{2})\cup(x_{3},x_{4})\cup\cdots$. If there exist $0<x_{1}<x_{2}<\dots<x_{k}$ such that $f$ has (weak) sign changes at $x_{1},x_{2},\dots,x_{k}$, we say that $f$ admits a finite (weak) sign change pattern.
@@
It is not true that for every smooth function $f\colon(0,+\infty)\to\mathbb{R}$ and every finite interval $a,b]\subset(0,+\infty)$, if $f$ takes both positive and negative values on $[a,b]$ then $f\big|_{[a,b]}$ admits a finite weak sign change pattern. However, it becomes true if “smooth” is replaced with “analytic” or “convex”.
@@
Let $h\colon(0,+\infty)\to\mathbb{R}$ be an integrable function. Suppose $\int_{0}^{+\infty}h(x)\mathrm{d}x=0$. If $h$ has exactly one weak sign change with sign pattern (0,0+)(0-,0+), and if $h$ is nonzero on a set of positive measure, then
@@
By the constraint that $h$ integrates to 0, we have for every $\alpha$,
@@
This concept appears naturally in moment problems, was advanced in [15], and has been extensively used in a variety of contexts, see [6, 7, 11, 17, 24, 29, 31, 32, 37]. Robust tools of this sort can be traced back to the classical work [25] of Karlin and Novikoff (with a slightly different focus on characterising certain cones of functions and their duals).
@@
(9) $X_{a,b}=a(\mathcal{E}-1)-b(\mathcal{E}^{\prime}-1).$
@@
(with the convention that for $a=0$, or $b=0$, the first, or the second case is empty, respectively, where we put $g_{a,b}\equiv 0$). In words, the distribution of $X_{a,b}$ is a mixture of two one-sided exponentials: $a\mathcal{E}-a+b$ with probability $\frac{a}{a+b}$ and $b\mathcal{E}^{\prime}-a+b$ with probability $\frac{b}{a+b}$.
@@
Let $X$ be a log-concave random variable with mean 0. There are unique $a,b\geq 0$ such that
@@
Let $\alpha=\mathbb{P}\left(X>0\right)$. Without loss of generality, we can assume that $\alpha\leq\frac{1}{2}$ (otherwise, we consider X-X instead of $X$). Since $\mathbb{E}X=0$, by Grünbaum’s theorem ([8, Lemma 2.2.6]), in fact, $\alpha\in[\frac{1}{e},\frac{1}{2}]$. Suppose that $a\geq b$. Then,
@@
Since the function $u\mapsto\frac{1}{e}\frac{e^{u}}{1+u}$ is continuous and strictly increasing on $0,{+\infty})$, it is bijective between $[0,1]$ and $\frac{1}{e},\frac{1}{2}]$. As a result, there is a unique value of the ratio $b/a$ for which $\mathbb{P}\left(X_{a,b}>0\right)=\alpha$. Furthermore, the mean of $X_{a,b}$ is 0, so
@@
[…]
@@
We proceed to show (12). By the homogeneity of $X/\mathbb{E}|X|$, we can assume that $\mathbb{E}|X|=1$. Let $f\colon\mathbb{R}\to[0,{+\infty})$ denote the density of $X$. For brevity, we shall write $g=g_{a,b}$ for the density of $X_{a,b}$ given above. These are log-concave probability densities, by construction $\log g$ is piecewise linear with maximum at $m=-(a-b)\leq 0$, enjoying the constraints
@@
Note that $f$ and $g$ on $m,{+\infty})$ can intersect at most twice, that is gfg-f changes sign at most twice because $\log f$ is concave and $\log g$ is linear. Similarly on $(-\infty,m]$. Because of the integral constraints on $0,{+\infty})$, Lemma 5 asserts that gfg-f needs to change its sign at least twice on $0,{+\infty})$, say at $0\leq x_{1}<x_{2}$, unless $f=g$ (identically). Then, gfg-f has the sign pattern (+,,+)(+,-,+) on $(0,{+\infty})$. Incidentally (it will be needed in the sequel), we also remark that $g-f\geq 0$ on $[m,0]$ as $\log g$ is linear on the entire half-line $m,{+\infty})$. We take $\lambda,\mu\in\mathbb{R}$ such that $\psi(x)-\lambda x-\mu$ vanishes at $x_{1},x_{2}$ ($\lambda,\mu$ are thus uniquely determined by the corresponding linear equations). Since $\psi$ is convex, $\psi(x)-\lambda x-\mu$ enjoys the sign pattern (0+,0,0+)(0+,0-,0+). Then,